# Asymmetric coupling of non-chaotic Rulkov neurons: fractal attractors, quasi-multistability, and final state sensitivity

### v1

@@ Abstract
Although neuron models have been well-studied for their rich dynamics and biological properties, limited research has been done on the complex geometries that emerge from the basins of attraction and basin boundaries of multistable neuron systems. In this paper, we investigate the geometrical properties of the strange attractors, four-dimensional basins, and fractal basin boundaries of an asymmetrically electrically coupled system of two identical non-chaotic Rulkov neurons. We discover a quasi-multistability in the system emerging from the existence of a chaotic spiking-bursting pseudo-attractor, and we classify and quantify the system’s basins of attraction, which are found to have complex fractal geometries. Using the method of uncertainty exponents, we also find that the system exhibits extreme final state sensitivity.
@@ I Introduction
The Rulkov maps [1, 2] are well-established to be simple discrete-time systems that accurately model the behavior of biological neurons. Although many continuous-time neuron models have been developed [3, 4, 5, 6, 7], the discrete-time Rulkov models have shown to have a computational advantage over these systems of nonlinear differential equations in modeling complex neuronal systems [8]. Like many other established neuron models [9, 10, 11, 12, 13, 14], Rulkov’s models, often called the non-chaotic and chaotic Rulkov maps [15], are slow-fast dynamical systems capable of modeling fast bursts of non-chaotic or chaotic spikes on top of slow oscillations. Although the chaotic Rulkov map has been well-studied in the literature [15, 16, 17, 18, 19, 20], this study involves non-chaotic Rulkov map, which has received less attention but also exhibits rich dynamical and geometrical properties. The most direct application of the non-chaotic Rulkov map is in modeling neuronal dynamics [1], but studies have also been conducted in using the non-chaotic Rulkov map for stability analysis [21], symbolic analysis [22], control of chaos [23], machine learning [24], information patterns [25], and digital watermarking [26]. In this paper, we explore an asymmetric electrical coupling of two identical non-chaotic Rulkov neurons and the emergence of a quasi-multistability, a strange spiking-bursting pseudo-attractor, and extreme final state sensitivity in the asymmetrically coupled Rulkov neuron system.
@@
When dynamical systems with attractors are analyzed, it is standard practice to detail their bifurcations, equilibria and eigenvalues, attractor dimensions, maximal Lyapunov exponents, and routes to chaos. However, less attention is often paid to a system’s basins of attraction, which are essential to determining multistability properties and the system’s use in practical applications [27]. Additionally, the fractalization of the boundaries between basins of attraction can result in final state sensitivity, where a small uncertainty in the initial state of a system results in a large uncertainty in the system’s eventual final state [28]. Even more extreme versions of final state sensitivity include Wada basins [29, 30, 31] and riddled basins [32, 33]. Despite these interesting and relevant geometrical properties, limited research has been done in the basins of attraction and basin boundaries of discrete-time neuronal systems. Some studies have been done with continuous-time neuron models, including multistability in a pair of neurons with time delays [34], heterogeneous multistability and riddled basins in an adaptive synapse-based neuron model [35], and multistability, fractal basin boundaries, and final state sensitivity in a network of theta neurons with time-varying excitability [36]. A few studies have also examined these properties in chaotic Rulkov neuron systems, such as multistability in a system of chaotic Rulkov neurons coupled via a chemical synapse [37], extreme multistability in a system of two chaotic Rulkov neurons with memristive electromagnetic induction [38], and basin stability in a system of two chaotic Rulkov neurons with a chemical synaptic and inner linking coupling [39]. However, to the best of our knowledge, neither the basins of attraction of a non-chaotic Rulkov neuron system nor the basin boundary and final state sensitivity properties of any discrete-time neuronal map have ever been studied before.
@@
The motivation of our study is therefore to analyze the complex geometrical properties that emerge from the basins of attraction and basin boundaries of a system of two asymmetrically coupled non-chaotic Rulkov neurons. In this system, we find a chaotic spiking-bursting pseudo-attractor and analyze its fractal geometry. Emerging from this chaotic pseudo-attractor and a separate non-chaotic spiking attractor is a kind of quasi-multistability and two basins of attraction. We classify and quantify these basins of attraction in multiple subsets of state space using the method of Sprott and Xiong [27], which has been a popular tool to analyze basins of attraction in recent years [40, 41, 42, 43, 44, 45, 46]. Then, we use uncertainty exponents [28, 47] to analyze the system’s extreme final state sensitivity and the fractalization of different subsets of the system’s basin boundaries using a Monte Carlo algorithm. This research provides novel insights into how multistability and fractal geometry can emerge from the basins of discrete-time neuronal systems, as well as how the slow-fast behavior of the non-chaotic Rulkov maps leads to complex and unexpected geometries in state space.
@@
This paper is organized as follows. In Section II, we detail our asymmetrically coupled non-chaotic Rulkov neuron model and the dynamics and apparent multistability that emerge from it. In Section III, we analyze the geometry of the chaotic spiking-bursting pseudo-attractor by computing its box-counting and Lyapunov dimensions. In Section IV, we overview the basin classification method and classify the intersection of the basins with a two-dimensional slice of state space (Section IV.1) and all of four-dimensional state space (Section IV.2). In Section V, we investigate different subsets of the system’s basin boundaries using uncertainty exponents and explore the existence of extreme final state sensitivity in the system. Finally, we summarize our results and give some future directions in Section VI.
@@ II The model and dynamics
The non-chaotic Rulkov map [1] is defined by the following iteration function  [^1] In the original paper that introduces the Rulkov map [1], the parameter $\sigma^{\prime}=\sigma+1$ is used, but we use the slightly modified form from [15].:
@@
Here, $x$ is the fast variable representing the voltage of the neuron, $y$ is the slow variable, and $\alpha$, $\sigma$, and $\mu$ are parameters. To make $y$ a slow variable, we need $0<\mu\ll 1$, so we choose the standard value of $\mu=0.001$. To couple Rulkov neurons together with an electrical current, we use a modified version of Eq. 1 that describes the dynamics of the $i$th coupled neuron $\mathbf{x}_{i}$ with the iteration function
@@
where $g^{e}_{ji}$ is the electrical coupling strength from the adjacent neuron $\mathbf{x}_{j}$ to $\mathbf{x}_{i}$ and the coefficients $\beta^{e}$ and $\sigma^{e}$ set the balance between the couplings of the neuron’s fast and slow variables [1].
In this paper, we examine a system of two asymmetrically electrically coupled non-chaotic Rulkov neurons, denoted by $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$. Since these neurons are coupled asymetrically, $g_{21}^{e}\neq g_{12}^{e}$, so we denote $g_{1}^{e}=g_{21}^{e}$ and $g_{2}^{e}=g_{12}^{e}$. Setting $\beta^{e}=\sigma^{e}=1$ for simplicity, Eqs. 4 and 5 read
@@
where $\mathbf{F}(\mathbf{X})$ can be written explicitly using Eqs. 3, 6, 7:
@@
The Jacobian matrix $J(\mathbf{X})$ of this system is quite complex due to the piecewise nature of the fast variable iteration function $f$ present in $\mathbf{F}(\mathbf{X})$ (see [49] for an explicit rendering of $J(\mathbf{X})$), but we can write it in a compact form by partitioning it using the method detailed in A:
@@
Figure 1: Graphs of the fast variable orbits $x_{1,k}$ and $x_{2,k}$ of two asymmetrically electrically coupled non-chaotic Rulkov neurons with parameters $\sigma_{1}=\sigma_{2}=-0.5$, $\alpha_{1}=\alpha_{2}=4.5$, and electrical coupling strengths $g^{e}_{1}=0.05$, $g^{e}_{2}=0.25$
For our system, we choose the asymmetrically coupled Rulkov neurons to be identical, with parameters $\sigma_{1}=\sigma_{2}=-0.5$ and $\alpha_{1}=\alpha_{2}=4.5$. This puts each of the individual neurons in the non-chaotic regular spiking domain. The coupling strengths we use are $g^{e}_{1}=0.05$ and $g^{e}_{2}=0.25$, which results in neuron $\mathbf{x}_{2}$ “feeling” the voltage difference between the two neurons more. To get a snapshot of the dynamics that can emerge from this model, we pick two slightly different initial conditions ($\mathbf{X}_{0}=(-0.54,-3.25,-1,-3.25)$ and $\mathbf{X}_{0}=(-0.56,-3.25,-1,-3.25)$) and plot the fast variable orbits of the two neurons in Fig. 1. In the first graph (Fig. 1(a)), the system exhibits non-chaotic, synchronized spiking with a negative maximal Lyapunov exponent $\lambda_{1}\approx-0.0057$. However, in the second graph (Fig. 1(b)), after changing the initial value of $x_{1}$ slightly, chaotic spiking-bursting occurs in both neurons, with a positive maximal Lyapunov exponent of $\lambda_{1}\approx 0.0207$. These two similar initial states immediately being attracted to two drastically different stable orbits, one chaotic and the other non-chaotic, implies the existence of multistability in this system.
@@
(c) Fig. 2(a) zoom
(d) Fig. 2(b) zoom
Figure 2: Graphs showing the attraction of two chaotic orbits to the system’s non-chaotic spiking attractor
@@
Figure 3: Projections of the four-dimensional non-chaotic spiking attractor of the asymmetrically coupled Rulkov neuron system
@@
Although Fig. 1 indicates that this system has two distinct attractors: a non-chaotic spiking attractor and a chaotic spiking-bursting attractor, further analysis of this system shows that this is not completely accurate. To see this, let us examine two initially chaotic orbits of this system. The first orbit we look at is the orbit from Fig. 1(b), which has the initial state $\mathbf{X}_{0}=(-0.56,-3.25,-1,-3.25)$. The second orbit is generated from the initial state $\mathbf{X}_{0}=(0.35,-3.25,-1.23,-3.25)$. In Figs 2(a) and 2(b), we graph 150,000 iterations of the fast variable orbits emerging from these two initial states. These graphs show that the initial states yield initially chaotic fast variable orbits, but the voltages eventually begin to spike periodically once they get attracted to the non-chaotic spiking attractor. This attraction is made especially clear in Figs. 2(c) and 2(d), which zoom into the region where the orbits transition from chaotic to non-chaotic. We propose that these dynamics occur because, during the period when the neurons are exhibiting chaotic spiking-bursting, their slow and fast variables happen to align by chance, which causes their voltages to latch onto each other’s spiking and propels the dynamics of the system to the non-chaotic spiking attractor. If this is true, then this system does not have a true chaotic attractor because, by chance, all orbits will eventually get attracted to the non-chaotic spiking attractor. This fact is also verified by our numerical simulations.
Therefore, we conclude that this system doesn’t exhibit true multistability, but rather, a quasi-multistability that is only observed in the short term. For this reason, we will treat the short-term chaotic dynamics exhibited by this system as occurring on a chaotic “pseudo-attractor” on which some orbits end up before being attracted to the non-chaotic spiking attractor.

### v3

# Asymmetric coupling of nonchaotic Rulkov neurons: Fractal attractors, quasimultistability, and final state sensitivity

@@ Abstract
Although neuron models have been well studied for their rich dynamics and biological properties, limited research has been done on the complex geometries that emerge from the basins of attraction and basin boundaries of multistable neuron systems. In this paper, we investigate the geometrical properties of the strange attractors, four-dimensional basins, and fractal basin boundaries of an asymmetrically electrically coupled system of two identical nonchaotic Rulkov neurons. We discover a quasimultistability in the system emerging from the existence of a chaotic spiking-bursting pseudo-attractor, and we classify and quantify the system’s basins of attraction, which are found to have complex fractal geometries. Using the method of uncertainty exponents, we also find that the system exhibits extreme final state sensitivity, which results in a dynamical uncertainty that could have important applications in neurobiology.
@@ I Introduction
Rulkov maps [1, 2] are well established to be simple discrete-time systems that accurately model the behavior of biological neurons. Although many continuous-time neuron models have been developed [3, 4, 5, 6, 7], the discrete-time Rulkov models have been shown to have a computational advantage over these systems of nonlinear differential equations in modeling complex neuronal systems [8]. Like many other established neuron models [6, 9, 10, 11, 12, 13], Rulkov’s models, often called the nonchaotic and chaotic Rulkov maps [14], are slow-fast dynamical systems capable of modeling fast bursts of nonchaotic or chaotic spikes on top of slow oscillations. Although the chaotic Rulkov map has been well studied [15, 16, 17, 18, 19], this study involves the nonchaotic Rulkov map, which has received less attention but also exhibits rich dynamical and geometrical properties. The most direct application of the nonchaotic Rulkov map is in modeling neuronal dynamics [1], but studies have also been conducted using the nonchaotic Rulkov map for stability analysis [20], symbolic analysis [21], control of chaos [22], machine learning [23], information patterns [24], and digital watermarking [25]. In this paper, we explore an asymmetric electrical coupling of two identical nonchaotic Rulkov neurons and the emergence of a quasimultistability, a strange spiking-bursting pseudo-attractor, and extreme final state sensitivity in the asymmetrically coupled Rulkov neuron system.
@@
When dynamical systems with attractors are analyzed, it is standard practice to detail their bifurcations, equilibria and eigenvalues, attractor dimensions, maximal Lyapunov exponents, and routes to chaos. However, less attention is often paid to a system’s basins of attraction, which are essential to determining multistability properties and the system’s use in practical applications [26]. Additionally, the fractalization of the boundaries between basins of attraction can result in final state sensitivity, where a small uncertainty in the initial state of a system results in a large uncertainty in the system’s eventual final state [27]. Even more extreme versions of final state sensitivity include Wada basins [28, 29, 30] and riddled basins [31, 32].
Despite these interesting and relevant geometrical properties, limited research has been done in the basins of attraction and basin boundaries of discrete-time neuronal systems. Some studies have been done with continuous-time neuron models, including multistability in a pair of neurons with time delays [33], heterogeneous multistability and riddled basins in an adaptive synapse-based neuron model [34], and multistability, fractal basin boundaries, and final state sensitivity in a network of theta neurons with time-varying excitability [35]. A few studies have also examined these properties in chaotic Rulkov neuron systems, such as multistability in a system of chaotic Rulkov neurons coupled via a chemical synapse [36], extreme multistability in a system of two chaotic Rulkov neurons with memristive electromagnetic induction [37], and basin stability in a system of two chaotic Rulkov neurons with a chemical synaptic and inner linking coupling [38]. However, as suggested by the name, the chaotic Rulkov map is designed to model the dynamics of chaotic spiking-bursting neurons [2], so it is not ideal for the purposes of examining basins of attraction and final-state sensitivity properties that arise from multistability. Instead, we are interested in a neuronal map that lends itself to coexisting qualitatively different behaviors, namely nonchaotic regular spiking and chaotic spiking-bursting neuronal dynamics, which the nonchaotic Rulkov map is suited for. However, the downside of using this map is that, while the fast variable of the chaotic Rulkov map is described by a simple iteration function, the nonchaotic Rulkov map’s fast variable is described by a piecewise function [see Eq. (2)], which makes analysis significantly more complex. For this reason, to the best of our knowledge, no prior studies have examined the basins of attraction of a nonchaotic Rulkov neuron system or the basin boundary and final state sensitivity properties of any discrete-time neuronal map.
@@
The motivation of our study is therefore to analyze the complex geometrical properties that emerge from the basins of attraction and basin boundaries of a system of two asymmetrically coupled nonchaotic Rulkov neurons. In this system, we find a chaotic spiking-bursting pseudo-attractor and analyze its fractal geometry. Emerging from this chaotic pseudo-attractor and a separate nonchaotic spiking attractor is a kind of quasimultistability and two basins of attraction. We classify and quantify these basins of attraction in multiple subsets of state space using the method of Sprott and Xiong [26], which has been a popular tool to analyze basins of attraction in recent years [39, 40, 41, 42, 43, 44, 45]. Then, we use uncertainty exponents [27, 46] to analyze the system’s extreme final state sensitivity and the fractalization of different subsets of the system’s basin boundaries using a Monte Carlo algorithm. This research provides insights into how multistability and fractal geometry can emerge from the basins of discrete-time neuronal systems, as well as how the slow-fast behavior of the nonchaotic Rulkov maps leads to complex and unexpected geometries in state space.
@@
This paper is organized as follows. In Sec. II, we detail our asymmetrically coupled nonchaotic Rulkov neuron model and the dynamics and apparent multistability that emerge from it. In Sec. III, we analyze the geometry of the chaotic spiking-bursting pseudo-attractor by computing its box-counting and Lyapunov dimensions. In Sec. IV, we overview the basin classification method and classify the intersection of the basins with a two-dimensional slice of state space (Sec. IV.1) and all of four-dimensional state space (Sec. IV.2). In Sec. V, we investigate different subsets of the system’s basin boundaries using uncertainty exponents and explore the existence of extreme final state sensitivity in the system. Finally, we summarize our results and give some future directions in Sec. VI.
@@ II The model and dynamics
The nonchaotic Rulkov map is a two-dimensional slow-fast map designed to be a simple model of the neuron. Unlike other neuron models composed of highly nonlinear systems of differential equations, the Rulkov map aims to simplify the complex dynamics of a neuron into a phenomenological model that can be used to better understand collective neuronal dynamics. In this way, a limitation of the Rulkov map is that it isn’t a completely accurate neuron model with variables and parameters that correspond with physical values with units. Instead, the Rulkov map mimics the behavior of biological neurons in an attempt to be used to aid in the understanding of more complex collective neuronal phenomena, the mechanisms of which can often be hidden behind the mathematical complexity of more biologically accurate neuron models. The main advantage of the Rulkov map is that, despite its simplicity, it is able to model a diverse variety of neuronal behaviors, including regular spiking, irregular spiking, nonchaotic bursts of spikes, and chaotic spiking-bursting [1]. Therefore, in this study, we use the Rulkov map to discover geometrical complexities and final state sensitivity of neuron systems in a phenomenological sense rather than as a prediction of direct experimental results.
The nonchaotic Rulkov map is defined by the following iteration function  [^1] In the original paper that introduces the Rulkov map [1], the parameter $\sigma^{\prime}=\sigma+1$ is used, but we use the slightly modified form from [14].:
@@
Here, $x$ is the fast variable representing the voltage of the neuron, $y$ is the slow variable, and $\alpha$, $\sigma$, and $\mu$ are parameters. To make $y$ a slow variable, we need $0<\mu\ll 1$, so we choose the standard value of $\mu=0.001$. To couple Rulkov neurons together with an electrical current, we use a modified version of Eq. (1) that describes the dynamics of the $i$th coupled neuron $\mathbf{x}_{i}$ with the iteration function
@@
where $g^{e}_{ji}$ is the electrical coupling strength from the adjacent neuron $\mathbf{x}_{j}$ to $\mathbf{x}_{i}$ and the coefficients $\beta^{e}$ and $\sigma^{e}$ set the balance between the couplings of the neuron’s fast and slow variables.
To motivate this choice of coupling where we add values proportional to the voltage difference to the slow variable $y$ and the effective value of $y$ in the fast map, we will briefly discuss the qualitative behavior of the nonchaotic Rulkov map  [^2] For a more in-depth discussion, see Refs. [1] and [49].. From Eq. (2), we can see that a higher value of the slow variable $y$ effectively raises the height of the fast variable iteration function $f$, which results in a quicker increase in $x$ before being reset by the third piece of $f$. Hence, higher values of $y$ result in shorter periodic orbits or more rapid spikes. It is easy to see from Eq. (1) that the parameter $\sigma$ is the value of $x$ required to keep $y$ constant. If $x<\sigma$, $y$ will slowly increase, and vice versa. Therefore, $\sigma$ is an “excitation parameter” for the Rulkov map: a higher value of $\sigma$ will result in $y$ to increase until the spikes are fast enough for the average value of $x$ to equal $\sigma$. From this, it is clear that $\sigma_{i,k}$ acts as a parameter representing gradual changes in dynamics resulting from a voltage difference between neurons, essentially resulting in a slow increase or decrease in the frequency of spikes depending on whether current is flowing in or out of the neuron. On the other hand, $\beta_{i,k}$ represents an immediate change in the dynamics of the neuron as it changes the effective value of $y$ for the slow variable $x$. In other words, if the voltage difference is positive and current is flowing into a neuron, the $\beta_{i,k}$ results in the slow variable $x$ “thinking” that the fast variable $y$ is larger, causing an immediate excitation. The effects of both coupling parameters are valuable to our study and have been observed in real biological neurons [1].
In this paper, we examine a system of two asymmetrically electrically coupled nonchaotic Rulkov neurons, denoted by $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$. Since these neurons are coupled asymetrically, $g_{21}^{e}\neq g_{12}^{e}$, so we denote $g_{1}^{e}=g_{21}^{e}$ and $g_{2}^{e}=g_{12}^{e}$. Setting $\beta^{e}=\sigma^{e}=1$ for simplicity, Eqs. (4) and (5) read
@@
where $\mathbf{F}(\mathbf{X})$ can be written explicitly using Eqs. (3), (6), and (7):
@@
The Jacobian matrix $J(\mathbf{X})$ of this system is quite complex due to the piecewise nature of the fast variable iteration function $f$ present in $\mathbf{F}(\mathbf{X})$ [see Ref. [49] for an explicit rendering of $J(\mathbf{X})$], but we can write it in a compact form by partitioning it using the method detailed in Appendix A:
@@
Figure 1: Graphs of the fast variable orbits $x_{1,k}$ and $x_{2,k}$ of two asymmetrically electrically coupled nonchaotic Rulkov neurons with parameters $\sigma_{1}=\sigma_{2}=-0.5$, $\alpha_{1}=\alpha_{2}=4.5$, and electrical coupling strengths $g^{e}_{1}=0.05$, $g^{e}_{2}=0.25$
For our system, we choose the asymmetrically coupled Rulkov neurons to be identical, with parameters $\sigma_{1}=\sigma_{2}=-0.5$ and $\alpha_{1}=\alpha_{2}=4.5$. This puts each of the individual neurons in the nonchaotic regular spiking domain. The coupling strengths we use are $g^{e}_{1}=0.05$ and $g^{e}_{2}=0.25$, which results in neuron $\mathbf{x}_{2}$ “feeling” the voltage difference between the two neurons more. To get a snapshot of the dynamics that can emerge from this model, we pick two slightly different initial conditions [$\mathbf{X}_{0}=(-0.54,-3.25,-1,-3.25)$ and $\mathbf{X}_{0}=(-0.56,-3.25,-1,-3.25)$] and plot the fast variable orbits of the two neurons in Fig. 1. In the first graph [Fig. 1(a)], the system exhibits nonchaotic, synchronized spiking with a negative maximal Lyapunov exponent $\lambda_{1}\approx-0.0057$. However, in the second graph [Fig. 1(b)], after changing the initial value of $x_{1}$ slightly, chaotic spiking-bursting occurs in both neurons, with a positive maximal Lyapunov exponent of $\lambda_{1}\approx 0.0207$. These two similar initial states immediately being attracted to two drastically different stable orbits, one chaotic and the other nonchaotic, implies the existence of multistability in this system.
@@
Figure 2: Graphs showing the attraction of two chaotic orbits to the system’s nonchaotic spiking attractor
@@
Figure 3: Projections of the four-dimensional nonchaotic spiking attractor of the asymmetrically coupled Rulkov neuron system
@@
Although Fig. 1 indicates that this system has two distinct attractors, a nonchaotic spiking attractor and a chaotic spiking-bursting attractor, further analysis of this system shows that this is not completely accurate. To see this, let us examine two initially chaotic orbits of this system. The first orbit we look at is the orbit from Fig. 1(b), which has the initial state $\mathbf{X}_{0}=(-0.56,-3.25,-1,-3.25)$. The second orbit is generated from the initial state $\mathbf{X}_{0}=(0.35,-3.25,-1.23,-3.25)$. In Fig. 2, we graph 150,000 iterations of the fast variable orbits emerging from these two initial states. These graphs show that the initial states yield initially chaotic fast variable orbits, but the voltages eventually begin to spike periodically once they get attracted to the nonchaotic spiking attractor. This attraction is made especially clear in Fig. 2 when we zoom into the region where the orbits transition from chaotic to nonchaotic. We propose that these dynamics occur because, during the period when the neurons are exhibiting chaotic spiking-bursting, their slow and fast variables happen to align by chance, which causes their voltages to latch onto each other’s spiking and propels the dynamics of the system to the nonchaotic spiking attractor. If this is true, then this system does not have a true chaotic attractor because, by chance, all orbits will eventually get attracted to the nonchaotic spiking attractor. This fact is also verified by our numerical simulations.
Therefore, we conclude that this asymmetrically coupled neuron system isn’t multistable. However, this system clearly does have something that is reminiscent of true multistability, with two coexisting qualitatively different behaviors that the orbits of different initial states exhibit. Indeed, we can utilize many of the techniques commonly used to analyze multistability in dynamical systems in the context of this neuron system. Therefore, we will call the phenomenon that this system exhibits “quasimultistability,” which represents a multistability that is observed only in the short term. In this system, numerical simulation suggests that quasimultistability is present for orbits of length on the order of $10^{4}$ (see Fig. 2). We suspect that this type of quasimultistability can occur only in coupled systems similar to this: a small number of coupled identical dynamical systems that exhibit periodic, nonchaotic behavior individually. In these specific systems, the coupling between individually nonchaotic systems can result in chaotic dynamics, and due to there being only a small number of identical systems, they can fall into synchronized periodic orbits in a finite amount of time. For our asymmetrically coupled neuron system, we will treat the short-term chaotic dynamics as occurring on a chaotic “pseudo-attractor” on which some orbits end up before being attracted to the nonchaotic spiking attractor. Despite the fact that the spiking attractor is the only true attractor of the system, we find that treating this chaotic spiking-bursting pseudo-attractor as an “attractor” as well results in rich geometries in state space emerging from this quasimultistability.